# A Perspective on Quantum Computing Applications in Quantum Chemistry using 25–100 Logical Qubits

### v1

@@ Abstract
The intersection of quantum computing and quantum chemistry represents a promising frontier for achieving quantum utility in domains of both scientific and societal relevance. Owing to the exponential growth of classical resource requirements for simulating quantum systems, quantum chemistry has long been recognized as a natural candidate for quantum computation. This perspective focuses on identifying scientifically meaningful use cases where early fault-tolerant quantum computers, which are considered to be equipped with approximately 25–100 logical qubits, could deliver tangible impact. We highlight near- to mid-term opportunities in algorithm and software design, discuss representative chemical problems suited for quantum acceleration, and propose strategic roadmaps and collaborative pathways for advancing practical quantum utility in quantum chemistry.
@@ I Introduction and Motivation
This year marks a significant milestone—roughly a century since the formulation of quantum mechanics fundamentally reshaped our understanding of matter at the atomic and molecular level [1, 2, 3]. The Schrödinger equation, introduced during this transformative period, provided the theoretical bedrock for quantum chemistry, enabling, in principle, the prediction of chemical properties and reactivity from first principles, famously demonstrated for the hydrogen molecule [4]. Over the subsequent 100 years, the field has witnessed impressive advances, with the development and application of sophisticated classical computational methods, such as Density Functional Theory (DFT) [5, 6, 7] and wave function-based approaches like Coupled Cluster (CC) theory [8, 9, 10], achieving remarkable success in explaining and predicting chemical phenomena [11, 12].
@@
Despite a century of progress and the power of established classical algorithms, substantial challenges remain. The inherent complexity of the quantum many-body problem, the very challenge laid barely by quantum mechanics itself, continues to limit the accuracy and applicability of classical methods, particularly in the treatment of (i) strongly correlated electronic systems (e.g., modeling complex catalytic sites like FeMoco [13]), (ii) the description of complex excited states crucial for photochemistry and materials science [14], (iii) the simulation of open quantum dynamics governing system-environment interactions [15], and (iv) the accurate prediction of weak interaction and thus transition state, which is critical to describe reaction kinetics [16]. These enduring limitations, stemming directly from the exponential growth of the Hilbert space with system size, motivate the exploration of fundamentally different computational paradigms. Quantum computational approaches, which leverage quantum phenomena directly, offer a potential pathway to supplement or even surpass classical techniques by tackling these intrinsically quantum problems more naturally, as envisioned early on by Feynman [17] and Lloyd [18], and extensively reviewed ever since [19, 20, 21].
This perspective focuses on identifying strategies, opportunities, and roadblocks relevant to performing quantum chemistry simulations using quantum hardware anticipated in the near-to-intermediate future—specifically, devices possessing approximately 25–100 logical qubits. This regime represents a crucial transitional phase, bridging the current era of noisy intermediate-scale quantum (NISQ) devices [22] and the long-term vision of large-scale, fully fault-tolerant quantum computing [23], known as Fault-Tolerant Application-Scale Quantum (FASQ) [24]. This intermediate scale presents a unique window of opportunity to potentially demonstrate impactful quantum computations, possibly achieving quantum utility for specific, well-chosen chemical problems. Success in this era will likely hinge on leveraging resource-aware algorithm design [25], developing robust hybrid quantum-classical workflows [26], and fostering interdisciplinary co-design between quantum algorithms, chemistry domain expertise, and hardware capabilities [27].
Realizing this opportunity necessitates a clear-eyed understanding of both the burgeoning capabilities, demonstrated by recent hardware milestones [28, 29, 30, 31], and the significant constraints of near-term fault-tolerant quantum hardware platforms. Before delving into specific quantum chemistry applications and the algorithms tailored for this 25–100 logical qubit regime, we first briefly overview the rapidly evolving landscape of quantum hardware development and further articulate the motivation for concentrating efforts within this specific computational resource range.
II The 25–100 Logical Qubit Regime: A Transitional Landscape
@@
Implementing a single logical qubit today demands thousands of physical qubits, depending on error rates and the chosen code, such as the surface code [23, 36, 37]. Nevertheless, improvements in coherence times, gate fidelities, and system integration across leading platforms—including superconducting qubits [38, 29], atom array [30], trapped ions [39, 40], and photonic architectures [41, 42]—suggest that processors comprising 25–100 logical qubits could become available within the next several years [43, 28, 44, 45].
@@
In comparison with long-term strategic directions, the 25–100 logical qubit regime marks a pivotal near-term threshold in the evolution of quantum computing applications in quantum chemistry (see Figure 1). Devices in this range may be the first to enable meaningful quantum chemistry simulations that are intractable on classical computers, such as accurately modeling complex molecules [13, 25]. However, the opportunity comes with substantial caveats: these early fault-tolerant systems will still face non-negligible logical error rates, limited coherence times relative to computation depth, and practical constraints on connectivity, measurement, and classical I/O [46].
Progress in this regime demands a re-evaluation of what constitutes “quantum utility” in chemistry [47]. It is not solely about outperforming classical solvers in compute time or accuracy for all problems, but rather about delivering new scientific insights into problems that are intrinsically quantum and difficult to treat or beyond classical methods—including strongly correlated electrons [10], quantum coherence in dynamics [48], and environmental interactions [15].
Development efforts are therefore increasingly centered around hybrid algorithms [26], embedding techniques [49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60], and variational methods [61] that operate with shallow circuits or make optimal use of limited logical qubit counts. These strategies aim to optimize the use of limited quantum resources while interfacing seamlessly with classical simulation frameworks.
@@
This regime is particularly well-suited for active space quantum simulations, where a judiciously chosen set of orbitals—often those associated with strong correlation or reactive behavior—is treated on a quantum computer, while classical components handle the weakly correlated environment. Techniques such as downfolding [62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72] and quantum embedding [49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60] provide viable pathways to construct effective Hamiltonians for such active spaces.
@@
In addition to ground-state energy estimation, quantum dynamics is emerging as an area where early utilities may arise [21, 48]. Simulating time-dependent processes, especially in open systems or photoinduced transformations, poses considerable challenges for classical methods due to memory bottlenecks and entanglement growth [73, 74]. Quantum processors may offer more efficient routes for handling such inherently quantum phenomena using algorithms like Trotterization or Qubitization [75, 76, 77].
Collectively, these considerations underscore the need for focused, resource-aware, and problem-specific approaches in the early fault-tolerant era of quantum computing. The following sections explore these directions in detail, with an emphasis on bridging algorithmic innovation, hardware development, and domain relevance.
@@ III Opportunities in Quantum Chemical Simulations
Quantum chemistry presents a set of grand challenges where quantum computers—even at the scale of 25–100 logical qubits—are expected to make meaningful contributions. Rather than aiming for a wholesale replacement of classical electronic structure methods, progress is anticipated through hybrid approaches that integrate quantum computing with classical techniques, high-performance computing (HPC), and artificial intelligence (AI). This synergy enables both targeted enhancement of computational workflows and the exploration of scientific regimes where classical models break down. In the following, we highlight several directions that offer promising opportunities for near- to mid-term demonstrations of quantum utility in quantum chemical simulations.
@@ III.1 Strong Correlation and Active Space Decomposition
Many chemically and industrially important systems—such as open-shell transition metal complexes [78, 79, 80] and -electron materials [81, 82, 83]—exhibit strong electronic correlation. This characteristic poses significant challenges for standard classical simulation methods. Density functional theory, for instance, while widely used in catalysis [84], faces limitations in quantitatively describing strong correlations, electron spin states in magnetic catalysts, and weak intermolecular forces, which can impact the accuracy of predicted reaction barriers. Single-reference wavefunction approaches also struggle with strong correlation [85, 86]. These limitations become particularly pronounced when high precision is essential, such as in modeling intricate catalytic mechanisms where quantitative accuracy is critical for mechanism identification and catalyst design, electronic excitations spanning valence and core levels [14, 87], and relativistic effects in heavy elements [88, 89, 90]. Capturing these properties often requires computationally demanding multi-reference methods, whose cost scales prohibitively [91, 92]. Consequently, these strongly correlated systems are prime candidates for quantum-accelerated solvers [20, 21, 93] and advanced embedding strategies [49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60], potentially overcoming classical barriers and handling complexities (like multiple active sites or interfaces in catalysis) currently beyond reach. Quantum algorithms applied within carefully chosen active spaces, derived via embedding or selection protocols, are already being explored for catalytic problems such as molecule-surface interactions [94, 95].
@@
A key insight facilitating these approaches is that electron correlation is often spatially localized within molecular fragments. This motivates fragment-based methods like the Localized Active Space (LAS) approach, which has emerged as a particularly promising framework [96, 97, 98, 99]. LAS constructs the total wavefunction as an antisymmetrized product of local active space wavefunctions defined on weakly entangled fragments. Each fragment is treated with high-level methods, while inter-fragment interactions are captured at a mean-field level. The LAS State Interaction (LASSI) refinement recovers spin symmetries by diagonalizing the full Hamiltonian within a basis of LAS-configured states. Notably, the inclusion of charge transfer (CT) configurations between fragments within the LASSI Hamiltonian can be crucial for achieving quantitative accuracy, for example, when calculating magnetic coupling constants in multinuclear complexes. The applicability and scalability of LAS/LASSI have been demonstrated through calculations on systems like Cr(III) dimers [100, 101, 98] and spin ladder in Fe3 compound [97], and large-scale LASSCF calculations are being applied to systems with over 1000 orbitals (e.g., Fe4S4, Cr2, NiFe2, and Ni2). LAS is part of a broader family of embedding techniques; related methods like Density Matrix Embedding Theory (DMET) [102] are also being combined with high-level solvers for large systems and specific applications such as core-level spectroscopy [103]. Together, these strategies provide a scalable foundation for integrating classical and quantum workflows, potentially targeting specific active sites identified by classical methods with high-accuracy quantum calculations [99, 104, 105].
@@
To further reduce complexity, downfolding methods can be employed to derive effective Hamiltonians that retain essential many-body physics within minimal active spaces [62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72]. Chemically motivated diagnostics—such as orbital occupation analysis and entanglement entropy—guide the selection of these reduced spaces [106], enabling the targeting of chemically significant regions using limited quantum resources.

### v2

# September 30, 2025 A Perspective on Quantum Computing Applications in Quantum Chemistry using 25–100 Logical Qubits

@@ Abstract
The intersection of quantum computing and quantum chemistry represents a promising frontier for achieving quantum utility in domains of both scientific and societal relevance. Owing to the exponential growth of classical resource requirements for simulating quantum systems, quantum chemistry has long been recognized as a natural candidate for quantum computation. This perspective focuses on identifying scientifically meaningful use cases where early fault-tolerant quantum computers, which are considered to be equipped with approximately 25–100 logical qubits, could deliver tangible impact. While recent advances in classical computing have pushed the boundaries of tractable simulations to unprecedented scales, this logical-qubit regime represents the first window where quantum devices can pursue qualitatively distinct strategies, such as polynomial-scaling phase estimation, direct simulation of quantum dynamics, and active-space embedding, that remain challenging for classical solvers, for instance, multireference charge-transfer and conical-intersection states central to photochemistry and materials design. We highlight near-term opportunities in algorithm and software design, discuss representative chemical problems suited for quantum acceleration, and propose strategic roadmaps and collaborative pathways for advancing practical quantum utility in quantum chemistry.
Contents
@@
- II The 25–100 logical-qubit regime: A Transitional Landscape
- III Opportunities in Quantum Chemical Simulations
- IV Algorithmic Innovations
- V Hybrid Integration Pathways
- VI Conclusion and Outlook
- VII Acknowledgements
- A Extended details on qumodes and cQED gate implementations
I Introduction and Motivation
@@
The year of 2025 marks a significant milestone—roughly a century since the formulation of quantum mechanics fundamentally reshaped our understanding of matter at the atomic and molecular level [1, 2, 3]. The Schrödinger equation, introduced during this transformative period, provided the theoretical bedrock for quantum chemistry, enabling, in principle, the prediction of chemical properties and reactivity from first principles, famously demonstrated for the hydrogen molecule [4]. Over the subsequent 100 years, the field has witnessed impressive advances, with the development and application of sophisticated classical computational methods, such as Density Functional Theory (DFT) [5, 6, 7] and wave function-based approaches like Coupled Cluster (CC) theory [8, 9, 10], achieving remarkable success in explaining and predicting chemical phenomena [11, 12].
@@
Despite a century of progress and the power of established classical algorithms, substantial challenges remain. The inherent complexity of the quantum many-body problem, the very challenge laid barely by quantum mechanics itself, continues to limit the accuracy and applicability of classical methods, particularly in the treatment of: (i) strongly correlated electronic systems, such as catalytic sites like FeMoco, where single-reference methods fail [13]; (ii) complex excited states crucial for photochemistry and materials science, such as conical intersections and charge-transfer states that underlie ultrafast photophysical processes [14]; (iii) open quantum dynamics that govern system–environment interactions [15]; and (iv) transition-state energetics and reaction barriers, where stretched bonds induce strong/static correlation beyond single-reference methods, and dispersion or environmental effects can secondarily modulate barrier heights [16]. These enduring limitations, stemming directly from the exponential growth of the Hilbert space with system size, motivate the exploration of fundamentally different computational paradigms. Quantum computational approaches, which leverage quantum phenomena directly, offer a potential pathway to supplement or even surpass classical techniques by tackling these intrinsically quantum problems more naturally, as envisioned early on by Feynman [17] and Lloyd [18], and extensively reviewed ever since [19, 20, 21].
Recent advances in classical computing have impressively extended the reach of exact and approximate electronic structure methods, including trillion-determinant full configuration interaction [22] and massively parallel relativistic density matrix renormalization group (DMRG) tailored coupled cluster approach [23]. These studies approach orbital ranges comparable to those accessible with 25–100 logical qubits under conventional mappings. Yet classical solvers remain subject to exponential barriers for general strongly correlated and dynamical problems. By contrast, even early fault-tolerant devices with 25–100 logical qubits can access qualitatively different algorithmic primitives, such as polynomial-scaling phase estimation, efficient Hamiltonian simulation, and block-encoded dynamics, that cannot be emulated efficiently at scale by classical algorithms.
Here, following recent usage in the community, we use the term quantum utility to mean reliable, validated quantum computations on domain-relevant tasks at scales beyond brute-force classical methods [24, 25, 26, 27]. This notion emphasizes practical, auditable performance on scientifically meaningful problems, reported with stated error bars and resource annotations (a standardized record of quantum resource requirements, including logical-qubit counts, circuit depth/-counts, Hamiltonian term norms, and shot budgets). It contrasts with quantum advantage, which often refers to worst-case or contrived separations, and with quantum dominance/supremacy, which targets formal separations on specially constructed tasks.
This perspective is motivated by the opportunity to translate these distinctive quantum capabilities into practical chemistry workflows. Achieving utility in the 25–100 logical-qubit regime will depend on resource-aware algorithm design [28], robust hybrid quantum-classical workflows [29], and interdisciplinary co-design across algorithms, chemistry, and hardware [30]. Realizing this opportunity also requires a clear-eyed view of both the burgeoning capabilities of emerging hardware [31, 32, 33, 34] and the practical constraints of near-term fault-tolerant platforms. Before turning to specific quantum chemistry applications and algorithms, we briefly overview this hardware landscape (Sec. II). Throughout, the 25–100 logical-qubit regime serves as a unifying lens for evaluating algorithmic strategies, benchmarking protocols, and collaborative pathways, with emphasis on resource-aware and auditable progress. The manuscript then proceeds from opportunities in quantum chemical simulations (Sec. III), to algorithmic innovations most relevant for this regime (Sec. IV), to hybrid integration pathways (Sec. V), and finally to the conclusion and outlook on co-design and collaboration models (Sec. VI).
II The 25–100 logical-qubit regime: A Transitional Landscape
In this section, we outline the evolving quantum hardware landscape and explain why the 25–100 logical-qubit regime marks a pivotal transitional window for quantum chemistry applications.
@@ II The 25–100 logical-qubit regime: A Transitional Landscape
Implementing a single logical qubit today demands many physical qubits, sometimes thousands, depending on error rates and the chosen code, such as the surface code [40, 41, 42]. Nevertheless, improvements in coherence times, gate fidelities, and system integration across leading platforms, including superconducting qubits [43, 32], atom array [33], trapped ions [44, 45], and photonic architectures [46, 47], suggest that processors comprising 25–100 logical qubits could plausibly become available on a 5–10 year horizon (see also a recent analysis of use cases from the NERSC workload, Ref. 48). This estimate reflects the scaling projections reported in recent national roadmaps [49, 50], IBM’s published timeline for error-corrected qubits [31], Google’s logical-qubit demonstrations [42, 32], and atom-array advances [33]. We stress that this horizon is a projection, not a guarantee, and depends critically on continued improvements in coherence, gate fidelity, and decoding throughput.
@@
In comparison with long-term strategic directions, the 25–100 logical-qubit regime marks a pivotal near-term threshold in the evolution of quantum computing applications in quantum chemistry (see Figure 1). Devices in this range may be the first to enable meaningful quantum chemistry simulations that are intractable on classical computers, such as accurately modeling complex molecules [13, 28]. Figure 1 situates these opportunities within a broader landscape, distinguishing near-term targets from longer-term aspirations. We emphasize that the long-term entries are illustrative and not the focus of this Perspective; our analysis centers on the 25–100 logical-qubit regime, corresponding to the near-term columns. In particular, strongly correlated systems (discussed in detail in Sec. III) are highlighted as compelling near-term candidates when treated within compact active spaces, rather than being deferred exclusively to the long-term horizon. However, the opportunity comes with substantial caveats: these early fault-tolerant systems will still face non-negligible logical error rates, limited coherence times relative to computation depth, and practical constraints on connectivity, measurement, and classical I/O [51].
Progress in this regime demands a re-evaluation of what constitutes “quantum utility” in chemistry [27]. It is not solely about outperforming classical solvers in compute time or accuracy for all problems, but rather about delivering new scientific insights into problems that are intrinsically quantum and difficult to treat or beyond classical methods, including strongly correlated electrons [10], quantum coherence in dynamics [52], and environmental interactions [15]. By scientific insights, we mean not just reproducing known quantities more efficiently, but accessing observables and regimes that classical simulations cannot capture reliably. Examples include uncovering mechanistic details of catalytic cycles that hinge on multireference transition states, resolving ultrafast photochemical processes via conical intersections, or quantifying coherence and dissipation effects in biomolecular dynamics. In these cases, even modest quantum devices may provide qualitative understanding, such as relative ordering of states, trends along reaction coordinates, or signatures of coherence, that extend beyond the reach of brute-force classical methods. black Development efforts are therefore increasingly centered around hybrid algorithms [29], embedding techniques [53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64], and variational methods [65] that operate with shallow circuits or make optimal use of limited logical qubit counts. These strategies aim to optimize the use of limited quantum resources while interfacing seamlessly with classical simulation frameworks.
@@
This regime is particularly well-suited for active space quantum simulations, where a judiciously chosen set of orbitals (often those associated with strong correlation or reactive behavior) is treated on a quantum computer, while classical components handle the weakly correlated environment. Techniques such as downfolding [66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76] and quantum embedding [53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64] provide viable pathways to construct effective Hamiltonians for such active spaces.
@@
In addition to ground-state energy estimation, quantum dynamics is emerging as an area where early utilities may arise [21, 52]. Simulating time-dependent processes, especially in open systems or photoinduced transformations, poses considerable challenges for classical methods due to memory bottlenecks and entanglement growth [77, 78]. Quantum processors, based on formal complexity bounds for Hamiltonian simulation [79, 80, 81], are projected to enable polynomial-scaling routes for handling such inherently quantum phenomena. Resource estimates in Table 2 show that even modest active spaces, such as Fe2S2 clusters or photochemical chromophores, map naturally to 25–40 logical qubits with depth – and shot budgets –, providing concrete targets for this regime.
Collectively, these considerations underscore the need for focused, resource-aware, and problem-specific approaches in the early fault-tolerant era of quantum computing. They set the stage for exploring chemical opportunities in Sec. III, where we examine problem classes and observables that align naturally with the 25–100 logical-qubit regime.
@@ III Opportunities in Quantum Chemical Simulations
In this section, we highlight problem classes and observables in quantum chemistry that align naturally with the 25–100 logical-qubit regime, emphasizing cases where quantum solvers can credibly complement or surpass classical approaches.
Quantum chemistry presents a set of grand challenges where quantum computers (even at the scale of 25–100 logical qubits) are expected to make meaningful contributions. Rather than aiming for a wholesale replacement of classical electronic structure methods, progress is anticipated through hybrid approaches that integrate quantum computing with classical techniques, high-performance computing (HPC), and artificial intelligence (AI). This synergy enables both targeted enhancement of computational workflows and the exploration of scientific regimes where classical models break down. In the following, we highlight several directions that offer promising opportunities for near-term demonstrations of quantum utility in quantum chemical simulations.
@@ III.1 Strong Correlation and Active Space Decomposition
Many chemically and industrially important systems, such as open-shell transition metal complexes [82, 83, 84] and -electron materials [85, 86, 87], exhibit strong electronic correlation. This characteristic poses significant challenges for standard classical simulation methods. Density functional theory, for instance, while widely used in catalysis [88], faces limitations in quantitatively describing strong correlations, electron spin states in magnetic catalysts, and noncovalent interactions (dispersion, hydrogen bonding), which can affect barrier predictions. Single-reference wave function approaches also struggle with strong correlation [89, 90]. These limitations become particularly pronounced when high precision is essential, such as in modeling intricate catalytic mechanisms where quantitative accuracy is critical for mechanism identification and catalyst design, electronic excitations spanning valence and core levels [14, 91], conical intersections governing photochemical branching, charge-transfer states in chromophores, and relativistic effects in heavy elements [92, 93, 94]. We note that these same considerations underlie the challenge of transition states: near the saddle point, partial bond cleavage induces strong/static (multi-reference) correlation and near-degeneracy that can render single-reference methods unreliable, motivating active-space, downfolded, or embedded treatments targeted along the reaction coordinate.
This exponential wall is clearly illustrated by the full configuration interaction (FCI) determinant growth,
where and are the numbers of - and -electrons, respectively. As can be seen, the growth increases combinatorially with the number of orbitals and electrons [22, 23]. Even modest active spaces with 20–30 orbitals already lead to billions of determinants, underscoring the limits of brute-force classical treatments. As mentioned in the Introduction, the orbital ranges addressed in recent distributed FCI and relativistic DMRG-tailored CC studies [22, 23] are comparable to those that can be mapped onto 25–100 logical qubits under conventional Jordan-Wigner or Bravyi-Kitaev fermion-to-qubit mappings, underscoring why this regime represents the first realistic window for quantum utility (see Sec. III.4 for representative systems and resource estimates).
Although the exponential wall ultimately limits all multi-reference methods, modern implementations have substantially extended their practical reach. Tensor network approaches such as DMRG (and its relativistic extensions), stochastic configuration interaction, and quantum Monte Carlo can now simulate thousands of orbitals on leadership-class HPC platforms. Their efficiency is, however, strongly system dependent: tensor networks are limited by entanglement growth and orbital ordering, quantum Monte Carlo by the fermion sign problem, and embedding approaches by fragment–bath partitioning and convergence control. These bottlenecks underscore both the remarkable reach of state-of-the-art classical solvers—often within carefully chosen active spaces—and the complementary potential of quantum algorithms to tackle the remaining hard cases.
Consequently, strongly correlated systems are prime candidates for quantum-accelerated solvers [20, 21, 95] and for advanced embedding strategies [53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64]. Embedding itself spans a spectrum: in some cases the correlation is spatially localized, motivating fragment-based treatments, while in other cases the correlation is delocalized across extended lattices, requiring systematic downfolding approaches (see Sec. IV.2 and Deep-Dive IV.2).
Electron correlation is not universally local: in strongly delocalized systems such as frustrated magnets or high- superconductors, correlation spans extended lattices and cannot be captured by fragment-based treatments. In such cases, downfolding techniques provide a more appropriate route to derive effective Hamiltonians that retain essential delocalized physics within compact forms. By contrast, in many molecular and bioinorganic systems, correlation effects are spatially concentrated within active sites or fragments. This motivates fragment-based strategies such as the Localized Active Space (LAS) approach, which has emerged as a promising framework for transition-metal complexes, molecular magnets, and bioinorganic clusters [96, 97, 98, 99].
@@
LAS constructs the total wavefunction as an antisymmetrized product of local active space wavefunctions defined on weakly entangled fragments. Each fragment is treated with high-level methods, while inter-fragment interactions are captured at a mean-field level. The LAS State Interaction (LASSI) refinement recovers spin symmetries by diagonalizing the full Hamiltonian within a basis of LAS-configured states. Notably, the inclusion of charge transfer (CT) configurations between fragments within the LASSI Hamiltonian can be crucial for achieving quantitative accuracy, for example, when calculating magnetic coupling constants in multinuclear complexes.
The applicability and scalability of LAS/LASSI have been demonstrated through calculations on systems like Cr(III) dimers [100, 101, 98] and spin ladder in Fe3 compound [97], and large-scale LASSCF calculations are being applied to systems with over 1000 orbitals (e.g., Fe4S4, Cr2, NiFe2, and Ni2). LAS is part of a broader family of embedding techniques; related methods like Density Matrix Embedding Theory (DMET) [102] are also being combined with high-level solvers for large systems and specific applications such as core-level spectroscopy [103]. Together, these strategies provide a promising foundation for integrating classical and quantum workflows, with practical scalability contingent on robust self-consistency [99, 104, 105]. While DMET offers appealing locality and parallelism, we note that single-shot and self-consistent variants can be sensitive to numerical fluctuations in the chemical-potential/correlation-potential update, particularly when the fragment solver is stochastic (e.g., on near-term quantum hardware). In practice, robust outer-loop strategies (e.g., DIIS/Anderson mixing, regularization of the correlation potential, and multi-moment matching beyond density) and reproducibility checks are required to attain stable convergence across fragments. At this point it is natural to ask why a quantum computer is needed at all if embedding reduces the problem to a compact active space. While many small active spaces (10–20 orbitals) remain tractable for advanced classical solvers, larger chemically relevant fragments (25–50+ orbitals) rapidly encounter exponential or entanglement-driven bottlenecks. We return to this point in Sec. III.4, where we connect active-space sizes to representative quantum resource estimates.
@@
To further reduce complexity, downfolding methods can be employed to derive effective Hamiltonians that retain essential many-body physics within minimal active spaces [66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76]. Chemically motivated diagnostics, such as orbital occupation analysis and entanglement entropy, can provide valuable guidance for selecting reduced spaces [106], but each comes with limitations: occupation numbers neglect two-particle correlation information, and entanglement measures are sensitive to the choice of orbital basis. Similarly, emerging AI-assisted strategies offer promising heuristics but remain an active area of research rather than a solved problem. Nevertheless, these approaches can inform the targeting of chemically significant regions, while highlighting the importance of cross-validation and reproducibility in active-space construction.